# Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models

### v3

@@ B.1 Benchmarks
Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
[…]
Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
Table 6: Results on the COCO subset of DetailCaps with more LVLMs.
@@
CHAIRi $=\frac{|\text{Hallucinated Sentences}|}{|\text{All Sentences}|}$
@@
is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al. (2023), Gemini-1.5-Pro Team et al. (2024), and GPT-4o OpenAI et al. (2024) for evaluation. It contains 4,870 images collected from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments. This benchmark primarily adopts the CAPTURE metric to evaluate caption quality. CAPTURE evaluates the alignment between generated and reference captions by calculating F1 scores using both hard and soft matching across entities ($F1_{obj}$), attributes ($F1_{attr}$), and relations ($F1_{rel}$). The final score is computed through a weighted aggregation of these components:
@@
Hal =$=\frac{\{\text{CHAIR}>0\}}{\{\text{All Caps}\}},$
@@ B.2 Evaluated LVLMs
To demonstrate the generalizability of the proposed CICD as a broadly applicable plug-and-play module, we evaluate its performance across six different LVLMs spanning three model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025). InstructBLIP and LLaVA-1.5 are both built upon Vicuna-7B Chiang et al. (2023) as their language backbone, while LLaVA-Next is based on LLaMA3. The Qwen-VL series, including Qwen2-VL-Instruct and Qwen2.5-VL-Instruct, is developed using the QwenLM framework. As for vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat utilize the pretrained CLIP model Radford et al. (2021), whereas Qwen2-VL-Instruct and Qwen2.5-VL-Instruct adopt a fine-tuned ViT Dosovitskiy et al. (2021). All models are implemented at the 8B parameter scale.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
[…]
LLaVA-1.5-13B Regular 58.8 17.0 73.4 82.94 81.46 78.49
CICD 44.4 12.9 76.0 87.92 85.68 80.99
INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
CICD 44.0 11.6 65.7 87.52 83.60 78.96
B.3 Baselines
We compare our method with several training-free baselines that specifically target language priors.
VCD Leng et al. (2024) introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination.
ICD Wang et al. (2024c) constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding.
IBD Zhu et al. (2024b) strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding.
DeFG Zhang et al. (2025) converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
Appendix C More Experiment Results
Figure 7: The correlation between similarity and performance. We conducted this experiment on the validation set of CHAIR. Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
C.1 Effectiveness on Larger LVLMs
@@
We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Tab.7. Our method continues to deliver significant performance improvements on a 13B model, indicating its strong generalizability. These results suggest that our approach scales well and remains effective across different model sizes.
C.2 Divergence Between Images
Contrastive decoding is a method that suppresses consistent features between two distributions while amplifying distinctive ones. To prevent the loss of visual information during contrastive decoding, we select an image with sufficient discrepancy to construct a contrasting distribution. Thus, while maintaining language prior consistency, the divergent visual information helps avoid erroneous debiasing caused by feature overlap. We use CLIP to retrieve images based on cosine similarity, and select the image with the lowest similarity as the contrastive image. We conduct experiments to explore the correlation between similarity and performance, as shown in Fig. 7.
We observe that the similarity between images exhibits no strong correlation with performance. Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements. As discussed in Sec.4.1, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships. This information exhibits limited overlap across real-world images. We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.8. Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs. In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
In Tab. 8, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency. We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion. Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
[…]
CHAIR(MaxLen=512)
Regular 54.6 16.4 72.6
w/o Image 54.6 15.5 76.6
Random Image 46.6 13.2 76.7
Retrieved Image 43.8 11.7 75.0
CHAIR(MaxLen=64)
Regular 24.4 8.9 56.6
w/o Image 25.4 7.6 61.9
Random Image 21.6 6.9 58.8
Retrieved Image 18.0 6.1 59.6
Table 8: This experiment was conducted by LLaVA-1.5. Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively. w/o Image represents the contrastive distribution that is constructed by removing the input image.
C.3 Threshold of Detrimental Priors
@@
While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency. We employ JS divergence to distinguish between essential and detrimental priors, where essential priors exhibit high cross-image consistency, reflected in lower JS divergence scores. We establish a threshold of $\gamma=-4$ in Eq.6, based on the statistical distribution of JS divergence scores shown in Fig.6.
@@
A comprehensive analysis is shown in Fig. 8. The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$. Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.
Appendix D Analysis of Visual Information Overlap
Fig. 2 presents the overlap ratio of image-related words between responses generated from negative contexts and those from regular decoding. Below, we detail the analysis procedure.
First, we apply several contrastive decoding methods, each constructing its own negative context, and use them to prompt LVLMs to answer user queries. For comparison, we obtain standard responses via regular decoding. We then use QwenVL-Max-0125 Bai et al. (2023b) to extract image-related words from the generated responses. Inspired by CoT Wei et al. (2022), we guide the model through a multi-stage extraction process: it first identifies all visually relevant words without access to the image, and then selects those strongly associated with the image content, which we define as relevant words.
Finally, we compute the overlap ratio between the relevant words generated from negative contexts and those obtained through regular decoding. This ratio quantifies how much essential visual content is preserved in responses generated from negative contexts, serving as an indicator of retained visual information. The prompt used to guide QwenVL-Max is shown in Fig. 9.
Based on this analysis, Fig. 2 shows that VCD and ICD retain over 30% of essential visual information, which is substantially weakened during contrastive decoding. In contrast, our method results in a negligible amount of residual information, effectively preventing visual information loss.
@@ Appendix D Analysis of Visual Information Overlap
Figure 8: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
Appendix E Case Study
We present case studies to illustrate the superior performance of our method in hallucination reduction, as shown in Fig. 10 and Fig. 11. The image captions are generated by LLaVA-1.5. Red highlights indicate hallucinated content, while green highlights denote content consistent with the image.
We can clearly observe that Regular Decoding produces a substantial amount of hallucinated content, often manifesting as content words such as objects, attributes, and relationships. In Fig. 10, for example, although the image depicts a skateboarding scene, the caption includes a reference to "spectators"—a detail not present in the image. Similarly, Fig. 11 shows an eating scene, yet the caption introduces a "long dining table," which is not actually visible. These hallucinations stem from LVLMs inheriting prior knowledge from LLMs, leading to biased descriptions in certain typical scenarios. In contrast, our method significantly reduces such hallucinations. While Fig. 11 still contains a hallucinated object (chips), this instance can be attributed to image ambiguity, where even human observers might plausibly infer such content.

### v5

@@ D.1 Benchmarks
CHAIRi $=\frac{|\text{Hallucinated Captions}|}{|\text{All Captions}|}$
@@
is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Wu et al. (2024), Gemini-1.5-Pro Team et al. (2024), and GPT-4o Hurst et al. (2024) for evaluation. It contains 4,870 images collected from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments. This benchmark primarily adopts the CAPTURE metric to evaluate caption quality. CAPTURE evaluates the alignment between generated and reference captions by calculating F1 scores using both hard and soft matching across entities ($F1_{obj}$), attributes ($F1_{attr}$), and relations ($F1_{rel}$). The final score is computed through a weighted aggregation of these components:
@@
Hal $=\frac{\{\text{CHAIR}>0\}}{\{\text{All Caps}\}},$
@@ D.2 Evaluated LVLMs
To demonstrate the generalizability of the proposed CICD as a broadly applicable plug-and-play module, we evaluate its performance across six different LVLMs spanning three model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024b), LLaVA-Next Liu et al. (2024c), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024a), and Qwen2.5-VL-Instruct Bai et al. (2025). InstructBLIP and LLaVA-1.5 are both built upon Vicuna-7B Chiang et al. (2023b) as their language backbone, while LLaVA-Next is based on LLaMA3. The Qwen-VL series, including Qwen2-VL-Instruct and Qwen2.5-VL-Instruct, is developed using the QwenLM framework. As for vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat utilize the pretrained CLIP model Radford et al. (2021), whereas Qwen2-VL-Instruct and Qwen2.5-VL-Instruct adopt a fine-tuned ViT Dosovitskiy et al. (2021). All models are implemented at the 8B parameter scale.
D.3 Results on CHAIR
The results in Table 7 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at a longer generation length. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
D.4 Results on AMBER
As illustrated in Table 8, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
D.5 Effectiveness on Larger LVLMs
@@
We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Table9. Our method continues to deliver significant performance improvements on a 13B model, indicating its strong generalizability. These results suggest that our approach scales well and remains effective across different model sizes.
D.6 Effects of $\gamma$
@@
While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency. We employ JS divergence to selectively suppress language priors, where priors exhibiting high cross-image consistency are retained, as reflected by lower JS divergence scores. We establish a threshold of $\gamma=-4$ in Eq. 2, based on the statistical distribution of JS divergence scores.
@@
A comprehensive analysis is shown in Figure 9. The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$. Another notable finding is that preserving necessary priors is beneficial, as it prevents the model’s generative capability from being compromised during the contrastive decoding process, thereby further reducing hallucinations.
@@
Figure 9: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$. $\gamma=$-$\infty$ is equivalent to removing both detrimental and essential priors.
@@
Appendix E Case Study
We present case studies to illustrate the superior performance of our method in hallucination reduction, as shown in Figure 11 and Figure 12. The image captions are generated by LLaVA-1.5. Red highlights indicate hallucinated content, while green highlights denote content consistent with the image.
We can clearly observe that Regular Decoding produces a substantial amount of hallucinated content, often manifesting as content words such as objects, attributes, and relationships. In Figure 11, for example, although the image depicts a skateboarding scene, the caption includes a reference to "spectators"—a detail not present in the image. Similarly, Figure 12 shows an eating scene, yet the caption introduces a "long dining table," which is not actually visible. These hallucinations stem from LVLMs inheriting prior knowledge from LLMs, leading to biased descriptions in certain typical scenarios. In contrast, our method significantly reduces such hallucinations.